OpenStudy (anonymous):

Previously, I asked for the derivative of -4cos (0.5x) and got 2sin(0.5x). However, I don't understand how we got the 2. I assume we're using the product rule.

OpenStudy (kira_yamato):

Chain rule. d/dx (-4cos (0.5x)) = -4 (-sin (0.5x)) (0.5) = 2 sin (0.5 x)

OpenStudy (campbell_st):

in your question ypu have -4cos(0.5x) thuse uses the chain rule let u = 0.5x du/dx = 0.5 y = -4cos(u) dy/du = 4sin(u) then using the chain rule its dy/dx = dy/du * du/dx = 4sin(u) * 0.5 = 2sin(u) resubstitute u = 0.5x gives dy/dx = 2sin(0.5x)

OpenStudy (anonymous):

Right, thanks, I have a better idea now. So we still retain -4. Do we always do that when we have a number in front of cos or sin? I originally thought that since it can be written as -4 * cos(0.5x) that -4 is derived into zero.

OpenStudy (kira_yamato):

I don't think so... Because it is a constant multiple so it's like y = k . f(x) ==> dy/dx = k . f'(x)

OpenStudy (kira_yamato):

In this case k=-4 and f(x) = cos(0.5x)

OpenStudy (anonymous):

Cool, thanks Kira Yamato.

OpenStudy (kira_yamato):

僕に感謝する必要はありません